Solve by the elimination method.
7x+9y=44
5x-9y=-92
What is the solution of the system? Select the correct choice and, if necessary, fill in the answer box to complete choice.
A. The solution is ____?
B. There are infiniely many solutions.
C. There is no solution

Answer by mananth(16949)

7x+9y=44 .............1
5x-9y=-92 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 1
7x+9y=44
5x-9y=-92
Add the two equations
12x=-48
/12
x=-4
plug value of x in (1)
7x+9y=44
-28+9y=44
9y=44+28
9y=72
y=8